A local political leader claims that 95% of the families in his area of responsibility 

were given “ayuda” during a one-month lockdown. Of a random sample of 200 families, 187 said they received relief or “ayuda”. Is this enough to affirm the 

leader’s claim? Use 𝑎 = 0.05.


1
Expert's answer
2022-06-01T11:13:44-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.95"

"H_a:p\not=0.95"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Evidence:

Based on the information provided, the significance level is "\alpha = 0.05

," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z|> 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{187/200-0.95}{\sqrt{\dfrac{0.95(1-0.95)}{200}}}=-0.9733"

Since it is observed that "|z| =0.9733<1.96= z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-0.9733)= 0.330404," and since "p=0.330404>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than 0.95, at the "\alpha = 0.05" significance level.